OpenStudy (babyslapmafro):

Please help me evaluate the following integral (click to see) using an appropriate substitution.

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\cos ^{4}(3t)\sin(3t)dt\]

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (babyslapmafro):

Thank you for doing that, but the answer provided in the attachment is different from the answer in the book.